OpenStudy (anonymous):

Can anyone please help me with this expression: (x-2)((L/2)-x-2)? I am trying to solve for L. I cannot seem to do it though, is there no solution for just L? (I am trying to isolate x in another expression by plugging in the solution of L, so that I can find x)

OpenStudy (anonymous):

Also, I am really not sure about what this actually means: "Use your answer to part (B) to write an expression (i.e. a function) of its exterior height x, and the wood length L." Does it mean that I am supposed to write a function of BOTH x and L, or a function of x, with L in it?

OpenStudy (badhi):

(x-2)((L/2)-x-2) is a expression, you cannot solve it (there's nothing to solve). can you post the complete question?

OpenStudy (anonymous):

It is sort of a series of questions, but basically, I am trying to solve this: "(A) Write an equation that expressed the required relationship between variables x,y, and the total length of the wood L. In Part 1, you had a linear equation of the form ax_by=50 where a and b were numbers. Now you will have a slightly more complicated equation involving x, y, and L, i.e. your equation will have three variables, x, y, and L in it.

OpenStudy (anonymous):

So I got 2x+2y=L for that.

OpenStudy (anonymous):

And then the second part is "(B) Solve your equation from part (A) in terms of x and L."

OpenStudy (anonymous):

I got y=(L/2)-x for that.

OpenStudy (anonymous):

And then this is the final part that I am stuck on: "(C) Use your answer to part (B) to write an expression (i.e. a function) of its exterior height x, and the wood length L. [The variable y will not appear in this equation]"

OpenStudy (anonymous):

And I am pretty sure I am supposed to solve for L, in that expression.

OpenStudy (anonymous):

I am basically trying to get L=[something] to plug into the original expression/function (x-2)((L/2)-x-2), to have only one variable, x, if that makes any sense.

OpenStudy (badhi):

it seems that you got (x-2)(L/2 -x -2) from (x-2)(y-2) what does this expression say physically?

OpenStudy (anonymous):

Well you see, I got (x-2)(L/2-x-2) from (x-2)(y-2) because I plugged in my solution for y=(L/2)-x into the y, as there is not supposed to be a variable y (x is the height, y is the width).

OpenStudy (badhi):

what I am asking is what does this expression mean. That is is this an area or something

OpenStudy (anonymous):

Oh, yes, it is the interior area of a picture frame.

OpenStudy (anonymous):

There was a first part, it is relatively simple as it gives you the length, but it is basically just finding the interior area of a hypothetical picture frame, where each part of the frame is an inch wide, and you form the frame in various ways depending on the Length.

OpenStudy (anonymous):

The reason why I am trying to solve this expression is so that I can plug in values of L that it gives me later on.

OpenStudy (anonymous):

Maybe I am not actually supposed to do it this way, I am not sure...

OpenStudy (badhi):

ok. So, Is the question asking the values of x, y so that you'd get the maximum area?

OpenStudy (anonymous):

Yeah, if it helps, the next question is "(D) Use your function to find the inside dimensions that will make this area a maximum. Now each of your inside dimensions will depend on L, i.e. they will be functions of L."

OpenStudy (badhi):

what you have to do is finding the length and the height of the frame in terms of "L" so that you'll get the maximum area for the frame. take the area \[A=(x-2)\left(\frac{L}{2} -x -2\right)= -x^2+\frac{L}{2}x -2\left(\frac{L}{2}-2\right)=-\left(x-\frac{L}{4}\right)^2+\frac{L^2}{16}-2\left(\frac{L}{2}-2\right)\] so, for A to become maximum, \[\left(x-\frac{L}{4}\right)^2\] has to become minimum which is 0. \[\begin{align*}\left(x-\frac{L}{4}\right)^2&=0\\ x&=\frac{L}{4}\end{align*}\] and \[y=\frac{L}{2}-x=\frac{L}{4}\]

OpenStudy (badhi):

If you know calculus, you can use \[\begin{align*} A(x)&=-x^2+\frac{L}{2}x-2\left(\frac{L}{2}-2\right)\\ \frac{dA}{dx}&=-2x+\frac{L}{2} \end{align*}\] when A is maximum \[\begin{align*}\frac{dA}{dx}&=0\\ -2x+\frac{L}{2}&=0 \\ x&=\frac{L}{4}\end{align*}\] you can find y as in the earlier post

OpenStudy (anonymous):

I am sorry, I get what you are saying, but in the beginning part, why is it \[-x^2+L/2-2(L/2-2)\]?

OpenStudy (anonymous):

Specifically, how did you get the 2(L/2-2)?

OpenStudy (badhi):

\[\begin{align*}(x-2)\left(\frac{L}{2}-x-2\right)&=(x-2)\left(-x+\left[\frac{L}{2}-2\right]\right)\\ &=-x^2+\left(\frac{L}{2}-2+2\right)x-2\left(\frac{L}{2}-2\right)\end{align*}\]

OpenStudy (anonymous):

OH! I see, I totally forgot keeping those in parentheses, thank you.

OpenStudy (badhi):

no problem. you're mostly welcome :)

OpenStudy (anonymous):

So how would I write these as a "function of L"? L(x)=[whatever makes it a maximum]?

OpenStudy (anonymous):

And to get that, I need to fully factor the expression, then set it to equal zero, then find each of the two possible inside dimensions?

OpenStudy (badhi):

function of L means write x in terms of L or x(L) not L(x)

OpenStudy (anonymous):

So how would I write that out? x(L)=?

OpenStudy (badhi):

we have already found it x(L) = L/4

OpenStudy (anonymous):

That is the only possible dimension? There is not another one?

OpenStudy (badhi):

it is the only answer. x=y=L/4

OpenStudy (anonymous):

I guess I am confused as to this "Now each of your inside dimensions will depend on L, i.e. they will be functions of L." There is only one inside dimension that will be a maximum?

OpenStudy (badhi):

not the dimension. It is the area that is maximized under the constrain of 2x+2y=L.

OpenStudy (anonymous):

What are the dimensions, then, of "(D) Use your function to find the inside dimensions that will make this area a maximum. Now each of your inside dimensions will depend on L, i.e. they will be functions of L."?

OpenStudy (anonymous):

One more thing, I am supposed to create a table of values that I get from plugging in whatever number into L, and I know that when L=50 the area must be 529.

OpenStudy (anonymous):

Just using reasoning. So I need some function dimensions of L that will agree with that solution.

OpenStudy (anonymous):

1. A picture framer has a piece of wood that measures one inch wide by 50 inches long. She would like to make a picture frame as shown below, and uses all 50 inches of wood she has. Notice that x and y in our sketch below are the outside dimensions of the frame. NOTE: In your paper, you will need to reproduce the figure and rewrite all of the questions as they appear in this paper. (A) Write an equation that expresses the required relationship between variables x, y, and the total length of the wood. (B) Solve your equation from part (A) for y in terms of x. (C) Write an expression (i.e. a function) that gives the interior area of the frame as a function of its exterior height x. (D) Find the values of x that make the interior area of the picture frame equal to zero. Explain why this is the case in the context of the picture frame problem. Draw the frames that correspond to these two extreme conditions. (E) Use your function to find the inside dimensions that will make this area a maximum. State also the maximum internal area of the picture frame. PART II 2. Now the picture framer would like to rework this problem for a piece of wood of arbitrary length L. [In PART I, L was 50 inches. Now L becomes a new variable so that we can generalize what we did in PART I for a piece of wood of any arbitrary length, L] (A) Write an equation that expresses the required relationship between variables x, y, and the total length of the wood L. In PART I, you had a linear equation of the form ax+by = 50 where a and b were numbers. Now you will have a slightly more complicated equation involving x, y and L, i.e. your equation will have three variables, x, y, and L in it. (B) Solve your equation from part (A) for y in terms of x and L. (C) Use your answer to part (B) to write an expression (i.e. a function) that gives the interior area of the frame as a function of its exterior height x, and the wood length L. [The variable y will not appear in this equation] (D) Use your function to find the inside dimensions that will make this area a maximum. Now each of your inside dimensions will depend on L, i.e. they will be functions of L. (E) Use your expressions for the inside dimensions to calculate the maximum area of each picture frame that can be built from each length of wood L in the table below. The first row is the problem you solved in PART I of this paper. Show at least one sample calculation for this part of the problem. Length L in Inches Area A in square inches 50 80 120

OpenStudy (anonymous):

It may actually be easier if you could just see the entire problem set. :P

OpenStudy (anonymous):

I REALLY appreciate this, by the way.

OpenStudy (badhi):

since you found out side dimensions i.e. x=y=L/4 the inside dimensions can be achieved from (x-2) and (y-2) and the area would be (x-2)(y-2)=((L/4)-2)^2. how did you get 529

OpenStudy (anonymous):

I just looked at everything I did so far, and I am honestly not sure how I got 529.

OpenStudy (anonymous):

Wouldn't 23+23=46 (as 50-4=46) and then 23x23=529 be the maximum internal area? I subtracted 4 because of the combined frame width of 4 inches, and we are only trying to find the maximum internal area.

OpenStudy (badhi):

you cannot substitute 4 from 50, take the lenght x and width y of the outside of the frame, 2x+2y=50 inner frame width would be (x-2) and (y-2). Then the length of the inner frame would be, 2(x-2)+2(y-2) not (x-2)+(y-2) then, 2(x-2)+2(y-2)=(2x+2y)-8=50-8 (not 50-4)